Please help me to solve this word problem:
Sally was able to paint a room in 4 hours and Joe did the same job in 6 hours, how long will it take the two of them to do it together? Thank you in anticipation of your answer.

Answer by jim_thompson5910(35256)

1/4 + 1/6 = 1/t

(1/4)(3/3) + (1/6)(2/2) = 1/t

(1*3)/(4*3) + (1*2)/(6*2) = 1/t

3/12 + 2/12 = 1/t

(3+2)/12 = 1/t

5/12 = 1/t

5t = 12*1

5t = 12

t = 12/5

t = 2 2/5

So it will take 2 2/5 hours, or 2.4 hours (which is 2 hours, 24 minutes) for them to do it together.